# Hodge diamonds of the Landau–Ginzburg orbifolds

### 26 Feb 2024 v2

@@ 3.1. Fixed loci of the ${\mathrm{GL}}_{f}$ elements
$\mathrm{Fix}(g):=\left\{(x_{1},\dots,x_{N})\in{\mathbb{C}}^{N}\ |\ g\cdot(x_{1},\dots,x_{N})=(x_{1},\dots,x_{N})\right\}.$ (12)
This is an eigenvalue $1$ subspace of ${\mathbb{C}}^{N}$ and therefore a linear subspace of ${\mathbb{C}}^{N}$. By $N_{g}:=\dim_{\mathbb{C}}\mathrm{Fix}(g)$ denote its dimension and by $f^{g}:=f|_{\mathrm{Fix}(g)}$ the restriction of $f$ to the fixed locus of $g$. For $g\in G_{f}^{d}$ this linear subspace is furthermore a span of a collection of standard basis vectors.
@@
$\mathrm{Fix}(h)=\{(x_{1},\dots,x_{N})\in{\mathbb{C}}^{N}~{}|~{}x_{j}=0,j\notin I_{h}\}.$
@@
Let $\widetilde{x}_{1},\dots,\widetilde{x}_{N}$ be the coordinates of ${\mathbb{C}}^{N}$ dual to the basis diagonalizing $g$. In this coordinates the polynomial $f^{g}$ is obtained by setting some of $\widetilde{x}_{\bullet}$ to zero. The proof follows now by the same argument as in Proposition 5 of [ET13]. ∎
Example 18.
If $f$ is of Fermat or loop type, then any $g\in{\mathrm{GL}}_{f}$, such that $g\neq\mathrm{id}$ satisfies $\mathrm{Fix}(g)=0$. If $f=x_{1}^{a_{1}}+x_{1}x_{2}^{a_{2}}+\dots x_{N-1}x_{N}^{a_{N}}$ is of chain type, then any $g\in{\mathrm{GL}}_{f}$, such that $g\neq\mathrm{id}$ satisfies $\mathrm{Fix}(g)=\{(x_{1},\dots,x_{p},0\dots,0)\in{\mathbb{C}}^{N}\ |\ x_{k}\in{\mathbb{C}}\}$ for some $p$ depending on $g$. The polynomial $f^{g}$ is of chain type again: $f^{g}=x_{1}^{a_{1}}+x_{1}x_{2}^{a_{2}}+\dots x_{p-1}x_{p}^{a_{p}}$.
@@ 3.2. Age of a ${\mathrm{GL}}_{f}$ element
Proposition 19.
1) For any $g\in{\mathrm{GL}}_{f}$ we have
$\mathrm{age}(g)+\mathrm{age}(g^{-1})=N-N_{g}=d_{g}.$ (14)
2) For a diagonalizable $g\in{\mathrm{GL}}_{f}$ we have ${\rm age}(g)=0$ if and only if $g=\mathrm{id}$.
3) We have $g\in{\mathrm{SL}}_{f}$ if and only if ${\rm age}(g)\in{\mathbb{N}}$.
3.3. Diagonal symmetries and a graph $\Gamma_{f}$
Proposition 20.
@@ 3.3. Diagonal symmetries and a graph $\Gamma_{f}$
We first show the statement for the arrows pointing at $k$. Having an arrow $j\to k$ means that $f$ has a monomial $x_{j}^{a_{j}}x_{k}$ as a summand with a nonzero coefficient. We have $g\cdot x_{k}\neq x_{k}$ and therefore the summand can only be preserved under the action of $g$ if $g\cdot x_{j}\neq x_{j}$. Having an oriented path $i\to j_{1}\to\dots\to j_{n}\to k$ we have by using the previous step that $g\cdot x_{j_{n}}\neq x_{j_{n}}$ and then $g\cdot x_{j_{a}}\neq x_{j_{a}}$ for all $a$. Hence, for $x_{i}$ too. ∎
@@
$G_{f}^{gr}:=\left\{(\lambda_{1},\dots,\lambda_{N})\in({\mathbb{C}}^{*})^{N}\,\left|\,\prod_{j=1}^{N}\lambda_{j}^{E_{1j}}=\dots=\prod_{j=1}^{N}\lambda_{j}^{E_{Nj}}=1\right\}\right.$ (15)
with $E_{ij}$ being the components of $E_{f}$. The group $G_{f}^{gr}$ is exactly the maximal group of diagonal symmetries of the difference $f-f_{add}$. In particular, every element of $G_{f}^{gr}$ preserves all graph monomials of $f$.
We have
$G_{f}^{d}\subseteq G_{f}^{gr}.$ (16)
and hence $G_{f}^{d}$ is a finite group.
An element ${\overline{g}=\dfrac{1}{r}(\alpha_{1},\dots,\alpha_{N})}$ belonging to $G_{f}^{gr}$ satisfies
$E_{f}\cdot\overline{g}\in{\mathbb{Z}}^{N}.$ (17)
This gives yet another characterization of the group $G_{f}^{gr}$
$G_{f}^{gr}\cong\left\{\overline{g}\in({\mathbb{Q}}/{\mathbb{Z}})^{N}\ |\ E_{f}\cdot\overline{g}\in{\mathbb{Z}}^{N}\right\}=E_{f}^{-1}{\mathbb{Z}}^{N}/{\mathbb{Z}}^{N}.$ (18)
It follows that every vector $\overline{g}$ giving a $G_{f}^{gr}$–element is a linear combination with integer coefficients of the columns of $E_{f}^{-1}$. Following the notation of Krawitz [K09] define $\overline{\rho}_{i}$ as the $i$-th column of $E_{f}^{-1}$
$E_{f}^{-1}=\left(\overline{\rho}_{1}|\dots|\overline{\rho}_{N}\right).$ (19)
@@
The elements $\rho_{k}$ generate $G_{f}^{gr}$ and $j_{f}=\rho_{1}\cdots\rho_{N}$. The columns of $E_{f}$ generate all relations on $\rho_{1},\dots,\rho_{N}$.
In particular, for $(E_{1k},\dots,E_{Nk})^{T}$ being a $k$–th column of $E_{f}$ we have in $G_{f}^{gr}$
@@
3.4. Diagonal symmetries of an invertible singularity
In [FJJS] for an invertible $f$ the authors gave the set ${\mathcal{S}}_{f}$ of all $N$–tuples $(s_{1},\dots,s_{N})$ such that every $g~{}\in~{}G_{f}^{d}~{}\backslash\{\mathrm{id}\}$ is written uniquely by
@@ 3.4. Diagonal symmetries of an invertible singularity
and $s_{k}=0$ if and only if $k\in I_{g}$. Due to Eq. (10) and Proposition 11 it’s enough to construct such set for Fermat, loop or chain type polynomials.
Proposition 21.
@@
  • (Fermat type): $11\leq s_{1}\leq a_{1}-$ 1
  • (loop type): $1\leq s_{k}\leq a_{k}$ and
    $s\neq(a_{1},1,a_{3},1,\dots,a_{N-1},1),\quad s\neq(1,a_{2},1,a_{4},1,\dots,a_{N})$ (20)
@@
  • (chain type): $s$ is of the form
@@
Corollary 22.
In the additive notation for the column $s^{T}=(s_{1},\dots,s_{N})^{T}$ we have
$\overline{g}=E_{f}^{-1}s^{T}.$
3.5. Diagonal symmetries of a quasihomogeneous singularity
For any quasihomogeneous singularity $f$ consider its graph decomposition as in Eq. (4). Up to the renumbering and rescaling of the variables we have
$f_{0}=x_{1}^{a_{1}}\quad\text{or}\quad f_{0}=x_{1}^{a_{1}}x_{2}+\dots+x_{K}^{a_{K}}x_{1},$ (22)
$f_{1}=x_{1}x_{K+1}^{b_{1}}+x_{K+1}x_{K+2}^{b_{2}}+\dots+x_{K+L-1}x_{K+L}^{b_{L}}$ (23)
@@ 3.5. Diagonal symmetries of a quasihomogeneous singularity
Any nontrivial $g\in G_{f_{0}}^{d}$ extends to an element $\widetilde{g}\in G_{f}^{gr}$. Moreover it follows that $\mathrm{Fix}(g)=0$ and also $\mathrm{Fix}(\widetilde{g})=0$ as long as $g\neq\mathrm{id}$. Similarly any element $h\in G_{f}^{gr}$ with $\mathrm{Fix}(h)=0$ acts nontrivially on $x_{1},\dots,x_{K}$ preserving $f_{0}$. Hence it defines $h_{0}\in G_{f_{0}}^{d}$ by the restriction.
At the same time any $h\in({\mathbb{C}}^{\ast})^{L}$ acting diagonally on $(x_{K+1},\dots,x_{K+L})$ preserving $f_{1}$ extends to an element of $G_{f}^{gr}$ assuming it to act trivially on $f_{0}$ and all other $f_{2},\dots,f_{p}$. One notes immediately that such elements $h$ are the elements of chain type polynomial symmetry group. Denote the group of all such elements by $G_{f_{1}}^{\circ}$.
@@
For a nontrivial element $g\in G_{f}^{gr}$ and its restriction $g_{0}\in G^{d}_{f_{0}}$ the extension $\widetilde{g}_{0}$ is not unique. However, having it fixed, we have by Proposition 20 that there is a unique set of $g_{k}\in G_{f_{k}}^{\circ}$ for $k=1,\dots,p$, s.t.
@@
We have that every $g_{k}$ acts non–trivially only on the variables of $f_{k}$ preserving all the variables of $f_{0}$ identically
@@
$|G_{f}^{gr}|=|G_{f_{0}}^{d}|\cdot|G_{f_{1}}^{gr}|\cdots|G_{f_{p}}^{gr}|.$
Associate to every $g_{0},g_{1},\dots,g_{p}$ an element $s_{0},s_{1},\dots,s_{p}$ as in Proposition 21. Composing them in one column $s$ we have
$\overline{g}=E_{f}^{-1}s^{T}.$
@@
Proposition 23.
For any $g\in G_{f}^{d}$, such that $\overline{g}=E_{f}^{-1}s$ we have
$\mathrm{age}(g)=(1,\dots,1)E_{f}^{-1}s^{T}.$
@@
We need to show that the components of $\overline{g}$ belong to $[0,1)$. This follows immediately from the equality $E_{f}\overline{g}=s$, the bounds on $s$ and the special form of the matrix $E_{f}$ (see Eq. (7)). ∎
3.6. Symmetries and the Calabi–Yau condition
@@
Let the matrix $E_{f}^{T}$ define a polynomial $f^{T}$. Namely, if for $f$ we have Eq (5), then
$f^{T}:=\sum_{k=1}^{N}c_{k}x_{1}^{E_{1k}}\cdots x_{N}^{E_{Nk}}.$
This polynomial does not necessarily define an isolated singularity. However it is quasihomogeneous again with some weights $q_{1}^{T},\dots,q_{N}^{T}$ by the same argument as in Proposition 10.
We call $f$ star–shaped if its graph $\Gamma_{f}$ consists of $N-1$ vertices all adjacent to one vertex. Namely, $f_{0}=x_{1}^{a_{1}}$, $p=N-1$ and $f_{i}=x_{1}x_{i+1}^{b_{i+1}}$. Such a polynomial satisfies the CY condition if and only if
@@
We will treat the star–shaped polynomials separately.
Proposition 24.
Let $f$ not being a star–shaped polynomial, satisfy the CY condition. Then the weights $q_{1}^{T},\dots,q_{N}^{T}$ are all positive.
@@
Let $E_{f}$ be written as in Eq. (7) and $A_{0}$ be a $K\times K$ loop type matrix as in Eq. (6). It’s immediate that $q_{K+1}^{T},\dots,q_{N}^{T}$ are positive.
For $i=1,\dots,K$ denote by ${\mathcal{A}}_{i}$ the sum of all $q_{j}^{T}$ with $j>K$, s.t. $j$–vertex of $\Gamma_{f}$ is adjacent to the $i$–th vertex.
Lemma 25.
@@
Let ${\mathcal{A}}_{i}\geq 1$ for some $i$. Let the vertices adjacent to the $i$–th vertex be labelled by $K+1,\dots,K+m$ contributing to $f$ with the monomials $x_{i}x_{K+1}^{b_{K+1}},\dots,x_{i}x_{K+m}^{b_{K+m}}$. Then
$q_{K+j}=\frac{1}{b_{K+j}}(1-q_{i})\quad\text{ and }\quad q_{K+j}^{T}\leq\frac{1}{b_{K+j}}.$
@@
If $S>1$ this gives $q_{i}\geq 1$ which contradicts the quasihomogeneity condition of $f$. If $S=1$ then $q_{K+1}+\dots+q_{K+m}=1-q_{i}$ and the CY condition can only hold if $f$ is a star–shaped CY polynomial. ∎
Let’s show that $q_{1}^{T}$ is positive. The proof for $q_{2}^{T},\dots,q_{N}^{T}$ is similar.
@@
$c_{i,i}$ $=\frac{a_{1}\cdots a_{K}}{a_{i}\det A},\qquad c_{i,i+r}=(-1)^{r}\frac{a_{1}\cdots a_{K}}{a_{i}\cdots a_{i+r}\det A},\quad 1\leq r\leq K-i,$
$c_{i,i-1}$ $=(-1)^{K-1}\frac{1}{\det A},\quad c_{i,i-r}=(-1)^{K-r}\frac{a_{i-r+1}\cdots a_{i-1}}{\det A},\quad 1\leq r\leq i-1.$
@@
$q_{i}^{T}=c_{1i}+\dots+c_{Ki}-c_{1i}{\mathcal{A}}_{1}-\dots-c_{Ki}{\mathcal{A}}_{K}.$
@@
$q_{1}^{T}$ $\geq c_{11}+\dots+c_{K1}+\frac{c_{11}}{1-c_{11}-\dots-c_{1K}}\left(\sum_{i,j=1}^{K}c_{ij}-1\right)$
$\qquad+\sum_{i=2}^{K}{\mathcal{A}}_{i}\left(-c_{i1}+c_{11}\frac{1-c_{i1}-\dots-c_{iK}}{1-c_{11}-\dots-c_{1K}}\right)$
$=\frac{(c_{11}+\dots+c_{K1})(1-c_{11}-\dots-c_{1K})+c_{11}\left(\sum_{ij=1}^{K}c_{ij}-1\right)}{1-c_{11}-\dots-c_{1K}}$
$\qquad+\sum_{i=2}^{K}{\mathcal{A}}_{i}\frac{c_{11}(1-c_{i1}-\dots-c_{iK})-c_{i1}(1-c_{11}-\dots-c_{1K})}{1-c_{11}-\dots-c_{1K}}.$
@@
$T_{r}$ $=a_{K-1}(\dots a_{r-2}(a_{r-1}(a_{r}-1)+1)-1)\dots+(-1)^{r-1}),$
$P_{K-1}$ $=a_{K-1},\quad P_{r}=a_{r}(T_{r+2}+P_{r+1}+(-1)^{r}),\quad r\leq K-2.$
@@
$(c_{11}+\dots+c_{K1})(1-c_{11}-\dots-c_{1K})+c_{11}\left(\sum_{ij=1}^{K}c_{ij}-1\right)=\frac{a_{K}}{\det A}(T_{3}+P_{2}-1).$
@@
$(1-c_{i1}-\dots-c_{iK})-c_{i1}(1-c_{11}-\dots-c_{1K})=\frac{a_{K}}{\det A}(T_{i}+(-1)^{i})\prod_{r=2}^{i-1}a_{i},\quad i\geq 3.$
@@
Proposition 26.
Let $f$ satisfy the CY condition. Then for any diagonalizable $g\in{\mathrm{GL}}_{f}$, such that $N_{g}=0$ we have
$\mathrm{age}(g)\geq\sum_{k=1}^{N}q_{k}.$ (24)
@@
The element $j_{f}$ is represented in the old and the new basis by the same diagonal matrix. The given element $g$ acts of $\widetilde{x}_{k}$ just by a rescaling. Therefore it’s enough to show the proposition for $g$ belonging to the maximal group of diagonal symmetries.
To prove the propositions for $g\in G_{f}^{d}$ it’s enough to prove the inequality for any $g\in G_{f}^{gr}$ with $N_{g}=0$ and $f$, such that the graph $\Gamma_{f}$ has only one connected component.

### 25 Mar 2024 v3

@@ 3.1 Fixed loci of the $\boldsymbol{{\mathrm{GL}}_{f}}$ elements
$\mathrm{Fix}(g):=\bigl{\{}(x_{1},\dots,x_{N})\in{\mathbb{C}}^{N}\mid g\cdot(x_{1},\dots,x_{N})=(x_{1},\dots,x_{N})\bigr{\}}.$
This is an eigenvalue $1$ subspace of ${\mathbb{C}}^{N}$ and therefore a linear subspace of ${\mathbb{C}}^{N}$. By $N_{g}:=\dim_{\mathbb{C}}\mathrm{Fix}(g)$ denote its dimension and by $f^{g}:=f|_{\mathrm{Fix}(g)}$ the restriction of $f$ to the fixed locus of $g$. For $g\in G_{f}^{d}$, this linear subspace is furthermore a span of a collection of standard basis vectors.
@@
$\mathrm{Fix}(h)=\bigl{\{}(x_{1},\dots,x_{N})\in{\mathbb{C}}^{N}\mid x_{j}=0,j\notin I_{h}\bigr{\}}.$
@@
Let $\widetilde{x}_{1},\dots,\widetilde{x}_{N}$ be the coordinates of ${\mathbb{C}}^{N}$ dual to the basis diagonalizing $g$. In this coordinates the polynomial $f^{g}$ is obtained by setting some of $\widetilde{x}_{\bullet}$ to zero. The proof follows now by the same argument as in [13, Proposition 5]. ∎
Example 3.6.
If $f$ is of Fermat or loop type, then any $g\in{\mathrm{GL}}_{f}^{d}$, such that $g\neq\mathrm{id}$ satisfies $\mathrm{Fix}(g)=0$. If $f=x_{1}^{a_{1}}+x_{1}x_{2}^{a_{2}}+\dots+x_{N-1}x_{N}^{a_{N}}$ is of chain type, then any $g\in{\mathrm{GL}}_{f}^{d}$, such that $g\neq\mathrm{id}$ satisfies $\mathrm{Fix}(g)=\bigl{\{}(x_{1},\dots,x_{p},0,\dots,0)\in{\mathbb{C}}^{N}\mid x_{k}\in{\mathbb{C}}\bigr{\}}$ for some $p$ depending on $g$. The polynomial $f^{g}$ is of chain type again: $f^{g}=x_{1}^{a_{1}}+x_{1}x_{2}^{a_{2}}+\dots+x_{p-1}x_{p}^{a_{p}}$.
@@ 3.2 Age of a $\boldsymbol{{\mathrm{GL}}_{f}}$ element
Proposition 3.7.
  • $(1)$
    For any $g\in{\mathrm{GL}}_{f}$ we have
    $\mathrm{age}(g)+\mathrm{age}\bigl{(}g^{-1}\bigr{)}=N-N_{g}=d_{g}.$ (3.2)
  • $(2)$
    For a diagonalizable $g\in{\mathrm{GL}}_{f}$ we have ${\rm age}(g)=0$ if and only if $g=\mathrm{id}$.
  • $(3)$
    We have $g\in{\mathrm{SL}}_{f}$ if and only if ${\rm age}(g)\in{\mathbb{N}}$.
3.3 Diagonal symmetries and a graph $\boldsymbol{\Gamma_{f}}$
Proposition 3.8.
@@ 3.3 Diagonal symmetries and a graph $\boldsymbol{\Gamma_{f}}$
We first show the statement for the arrows pointing at $k$. Having an arrow $j\to k$ means that $f$ has a monomial $x_{j}^{a_{j}}x_{k}$ as a summand with a nonzero coefficient. We have $g\cdot x_{k}\neq x_{k}$ and therefore the summand can only be preserved under the action of $g$ if $g\cdot x_{j}\neq x_{j}$. Having an oriented path $i\to j_{1}\to\cdots\to j_{n}\to k$ we have by using the previous step that $g\cdot x_{j_{n}}\neq x_{j_{n}}$ and then $g\cdot x_{j_{a}}\neq x_{j_{a}}$ for all $a$. Hence, for $x_{i}$ too. ∎
@@
$G_{f}^{\rm gr}:=\Biggl{\{}(\lambda_{1},\dots,\lambda_{N})\in({\mathbb{C}}^{*})^{N}\,\Biggl{|}\,\prod_{j=1}^{N}\lambda_{j}^{E_{1j}}=\dots=\prod_{j=1}^{N}\lambda_{j}^{E_{Nj}}=1\Biggr{\}}$
with $E_{ij}$ being the components of $E_{f}$. The group $G_{f}^{\rm gr}$ is exactly the maximal group of diagonal symmetries of the difference $f-f_{\rm add}$. In particular, every element of $G_{f}^{\rm gr}$ preserves all graph monomials of $f$.
We have $G_{f}^{d}\subseteq G_{f}^{\rm gr}$ and hence $G_{f}^{d}$ is a finite group. An element ${\overline{g}=\frac{1}{r}(\alpha_{1},\dots,\alpha_{N})}$ belonging to $G_{f}^{\rm gr}$ satisfies $E_{f}\cdot\overline{g}\in{\mathbb{Z}}^{N}$. This gives yet another characterization of the group $G_{f}^{\rm gr}$
$G_{f}^{\rm gr}\cong\bigl{\{}\overline{g}\in({\mathbb{Q}}/{\mathbb{Z}})^{N}\mid E_{f}\cdot\overline{g}\in{\mathbb{Z}}^{N}\bigr{\}}=E_{f}^{-1}{\mathbb{Z}}^{N}/{\mathbb{Z}}^{N}.$
It follows that every vector $\overline{g}$ giving a $G_{f}^{\rm gr}$-element is a linear combination with integer coefficients of the columns of $E_{f}^{-1}$. Following the notation of Krawitz [21] define $\overline{\rho}_{i}$ as the $i$-th column of $E_{f}^{-1}$
$E_{f}^{-1}=(\overline{\rho}_{1}|\dots|\overline{\rho}_{N}).$
@@
The elements $\rho_{k}$ generate $G_{f}^{\rm gr}$ and $j_{f}=\rho_{1}\cdots\rho_{N}$. The columns of $E_{f}$ generate all relations on $\rho_{1},\dots,\rho_{N}$.
In particular, for $(E_{1k},\dots,E_{Nk})^{\mathsf{T}}$ being a $k$-th column of $E_{f}$ we have in $G_{f}^{\rm gr}$
@@
3.4 Diagonal symmetries of an invertible singularity
In [14] for an invertible $f$, the authors gave the set ${\mathcal{S}}_{f}$ of all $N$-tuples $(s_{1},\dots,s_{N})$ such that every $g~{}\in~{}G_{f}^{d}~{}\backslash\{\mathrm{id}\}$ is written uniquely by
@@ 3.4 Diagonal symmetries of an invertible singularity
and $s_{k}=0$ if and only if $k\in I_{g}$. Due to equation (3.1) and Proposition 2.10, it is enough to construct such set for Fermat, loop or chain type polynomials.
Proposition 3.9.
@@
  • ((((Fermat type)))): $11\leq s_{1}\leq a_{1}-$ 1
  • ((((loop type)))): $1\leq s_{k}\leq a_{k}$ and
    $s\neq(a_{1},1,a_{3},1,\dots,a_{N-1},1),\qquad s\neq(1,a_{2},1,a_{4},1,\dots,a_{N})$
@@
  • ((((chain type)))): $s$ is of the form
@@
Corollary 3.10.
In the additive notation for the column ${s^{\mathsf{T}}=(s_{1},\dots,s_{N})^{\mathsf{T}}}$, we have ${\overline{g}=E_{f}^{-1}s^{\mathsf{T}}}$.
3.5 Diagonal symmetries of a quasihomogeneous singularity
For any quasihomogeneous singularity $f$, consider its graph decomposition as in equation (2.2). Up to the renumbering and rescaling of the variables, we have
$f_{0}=x_{1}^{a_{1}}\qquad\text{or}\qquad f_{0}=x_{1}^{a_{1}}x_{2}+\dots+x_{K}^{a_{K}}x_{1},$
$f_{1}=x_{1}x_{K+1}^{b_{1}}+x_{K+1}x_{K+2}^{b_{2}}+\dots+x_{K+L-1}x_{K+L}^{b_{L}}$
@@ 3.5 Diagonal symmetries of a quasihomogeneous singularity
Any nontrivial $g\in G_{f_{0}}^{d}$ extends to an element $\widetilde{g}\in G_{f}^{\rm gr}$. Moreover it follows that $\mathrm{Fix}(g)=0$ and also $\mathrm{Fix}(\widetilde{g})=0$ as long as $g\neq\mathrm{id}$. Similarly any element $h\in G_{f}^{\rm gr}$ with $\mathrm{Fix}(h)=0$ acts nontrivially on $x_{1},\dots,x_{K}$ preserving $f_{0}$. Hence it defines $h_{0}\in G_{f_{0}}^{d}$ by the restriction.
At the same time any $h\in({\mathbb{C}}^{\ast})^{L}$ acting diagonally on $(x_{K+1},\dots,x_{K+L})$ preserving $f_{1}$ extends to an element of $G_{f}^{\rm gr}$ assuming it to act trivially on $f_{0}$ and all other $f_{2},\dots,f_{p}$. One notes immediately that such elements $h$ are the elements of chain type polynomial symmetry group. Denote the group of all such elements by $G_{f_{1}}^{\circ}$.
@@
For a nontrivial element $g\in G_{f}^{\rm gr}$ and its restriction $g_{0}\in G^{d}_{f_{0}}$, the extension $\widetilde{g}_{0}$ is not unique. However, having it fixed, we have by Proposition 3.8 that there is a unique set of $g_{k}\in G_{f_{k}}^{\circ}$ for $k=1,\dots,p$, s.t.
@@
We have that every $g_{k}$ acts non-trivially only on the variables of $f_{k}$ preserving all the variables of $f_{0}$ identically.
@@
$|G_{f}^{\rm gr}|=|G_{f_{0}}^{d}|\cdot|G_{f_{1}}^{\rm gr}|\cdots|G_{f_{p}}^{\rm gr}|.$
Associate to every $g_{0},g_{1},\dots,g_{p}$ an element $s_{0},s_{1},\dots,s_{p}$ as in Proposition 3.9. Composing them in one column $s$, we have
$\overline{g}=E_{f}^{-1}s^{\mathsf{T}}.$
@@
Proposition 3.11.
For any $g\in G_{f}^{d}$ such that $\overline{g}=E_{f}^{-1}s$, we have
$\mathrm{age}(g)=(1,\dots,1)E_{f}^{-1}s^{\mathsf{T}}.$
@@
We need to show that the components of $\overline{g}$ belong to $[0,1)$. This follows immediately from the equality $E_{f}\overline{g}=s$, the bounds on $s$ and the special form of the matrix $E_{f}$ (see equation (2.5)). ∎
3.6 Symmetries and the Calabi–Yau condition
@@
Let the matrix $E_{f}^{\mathsf{T}}$ define a polynomial $f^{\mathsf{T}}$. Namely, if for $f$ we have (2.3), then
$f^{\mathsf{T}}:=\sum_{k=1}^{N}c_{k}x_{1}^{E_{1k}}\cdots x_{N}^{E_{Nk}}.$
This polynomial does not necessarily define an isolated singularity. However, it is quasihomogeneous again with some weights $q_{1}^{\mathsf{T}},\dots,q_{N}^{\mathsf{T}}$ by the same argument as in Proposition 2.9.
We call $f$ star-shaped if its graph $\Gamma_{f}$ consists of $N-1$ vertices all adjacent to one vertex. Namely, $f_{0}=x_{1}^{a_{1}}$, $p=N-1$ and $f_{i}=x_{1}x_{i+1}^{b_{i+1}}$. Such a polynomial satisfies the CY condition if and only if
@@
We will treat the star-shaped polynomials separately.
Proposition 3.12.
Let $f$ not being a star-shaped polynomial, satisfy the CY condition. Then the weights $q_{1}^{\mathsf{T}},\dots,q_{N}^{\mathsf{T}}$ are all positive.
@@
Let $E_{f}$ be written as in equation (2.5) and $A_{0}$ be a $K\times K$ loop type matrix as in equation (2.4). It is immediate that $q_{K+1}^{\mathsf{T}},\dots,q_{N}^{\mathsf{T}}$ are positive.
For $i=1,\dots,K$, denote by ${\mathcal{A}}_{i}$ the sum of all $q_{j}^{\mathsf{T}}$ with $j>K$, s.t. $j$-vertex of $\Gamma_{f}$ is adjacent to the $i$-th vertex.
Lemma 3.13.
@@
Let ${\mathcal{A}}_{i}\geq 1$ for some $i$. Let the vertices adjacent to the $i$-th vertex be labelled by $K+1,\dots,K+m$ contributing to $f$ with the monomials $x_{i}x_{K+1}^{b_{K+1}},\dots,x_{i}x_{K+m}^{b_{K+m}}$. Then
$q_{K+j}=\frac{1}{b_{K+j}}(1-q_{i})\qquad\text{and}\qquad q_{K+j}^{\mathsf{T}}\leq\frac{1}{b_{K+j}}.$
@@
If $S>1$, this gives $q_{i}\geq 1$ which contradicts the quasihomogeneity condition of $f$. If $S=1$, then $q_{K+1}+\dots+q_{K+m}=1-q_{i}$ and the CY condition can only hold if $f$ is a star-shaped CY polynomial. ∎
Let us show that $q_{1}^{\mathsf{T}}$ is positive. The proof for $q_{2}^{\mathsf{T}},\dots,q_{N}^{\mathsf{T}}$ is similar.
@@
$c_{i,i}=\frac{a_{1}\cdots a_{K}}{a_{i}\det A},\qquad c_{i,i+r}=(-1)^{r}\frac{a_{1}\cdots a_{K}}{a_{i}\cdots a_{i+r}\det A},\qquad 1\leq r\leq K-i,$
$c_{i,i-1}=(-1)^{K-1}\frac{1}{\det A},\qquad c_{i,i-r}=(-1)^{K-r}\frac{a_{i-r+1}\cdots a_{i-1}}{\det A},\qquad 1\leq r\leq i-1.$
@@
$q_{i}^{\mathsf{T}}=c_{1i}+\dots+c_{Ki}-c_{1i}{\mathcal{A}}_{1}-\dots-c_{Ki}{\mathcal{A}}_{K}$
@@
$q_{1}^{\mathsf{T}}\geq{}$ $c_{11}+\dots+c_{K1}+\frac{c_{11}}{1-c_{11}-\dots-c_{1K}}\left(\sum_{i,j=1}^{K}c_{ij}-1\right)$
$+\sum_{i=2}^{K}{\mathcal{A}}_{i}\left(-c_{i1}+c_{11}\frac{1-c_{i1}-\dots-c_{iK}}{1-c_{11}-\dots-c_{1K}}\right)$
[…]
$+\sum_{i=2}^{K}{\mathcal{A}}_{i}\frac{c_{11}(1-c_{i1}-\dots-c_{iK})-c_{i1}(1-c_{11}-\dots-c_{1K})}{1-c_{11}-\dots-c_{1K}}.$
@@
$T_{r}=a_{K-1}(\cdots a_{r-2}(a_{r-1}(a_{r}-1)+1)-1)\cdots+(-1)^{r-1},$
$P_{K-1}=a_{K-1},\qquad P_{r}=a_{r}(T_{r+2}+P_{r+1}+(-1)^{r}),\qquad r\leq K-2.$
@@
$(c_{11}+\dots+c_{K1})(1-c_{11}-\dots-c_{1K})+c_{11}\left(\sum_{ij=1}^{K}c_{ij}-1\right)=\frac{a_{K}}{\det A}(T_{3}+P_{2}-1)$
@@
$(1-c_{i1}-\dots-c_{iK})-c_{i1}(1-c_{11}-\dots-c_{1K})=\frac{a_{K}}{\det A}\left(T_{i}+(-1)^{i}\right)\prod_{r=2}^{i-1}a_{i},\qquad i\geq 3.$
@@
Proposition 3.14.
Let $f$ satisfy the CY condition. Then for any diagonalizable $g\in{\mathrm{GL}}_{f}$ such that $N_{g}=0$, we have
$\mathrm{age}(g)\geq\sum_{k=1}^{N}q_{k}.$
@@
The element $j_{f}$ is represented in the old and the new basis by the same diagonal matrix. The given element $g$ acts of $\widetilde{x}_{k}$ just by a rescaling. Therefore it is enough to show the proposition for $g$ belonging to the maximal group of diagonal symmetries.
To prove the propositions for $g\in G_{f}^{d}$ it is enough to prove the inequality for any $g\in G_{f}^{\rm gr}$ with $N_{g}=0$ and $f$, such that the graph $\Gamma_{f}$ has only one connected component.